40 Chapter 1 5. The electric bill for the period of December 15, 20X5, through January 10, 20X6 totaled $320. It has not been recorded and will be paid on January 15, 20x6. Required: For each situation, prepare the adjusting entry to record the proper expense for addition, prepare the journal entry to write off the two uncollectible accounts. Problem 14 Below is a list of the 20X3 activities/transactions for Emilys Eatery: 1. A fire insurance policy for July 1, 20X3, through June 30, 20X4, was purchased on May 15, 20X3. The premium paid totaled $24,000 and was recorded as prepaid insurance 2. The operation purchased a new cash register costing $12,000 on September 1, 20x3, and recorded it in the equipment account. The cash register is expected to have a use- ful life of seven years and a salvage value of $1,000. No depreciation has been recorded Emilys uses the straight-line method of depreciation. Ad 3. The employees were paid for their work through December 26. They worked 300 4. Sales for the year totaled $800,000. The allowance for doubtful accounts has a 5. The electric bill for the period of December 5, 20X3, through January Required hours for the period of December 27-31, 20X3, and will be paid on January 10, 20X4. The average hourly wage is $6.00 1. Decem- ber 31 balance of $1,000. The allowance should be adjusted at year-end to /h perce sales for the year. 2. 3. $620. It has not been recorded and will be paid on January 10, 20X4. 4. For each situation, prepare the adjusting entnu to for 20X3.Based on the information provided, prepare a simple income statement based on cash activity (cash accounting). 2. Based on the information provided, prepare a simple income statement based on rev- enue earned and actual expenses incurred (accrual accounting) 3. What is the amount of cash on hand at the end of the first month? Problem 13 Below is a list of the activities/transactions for Fees Fountain: A fire insurance policy for January 1, 20x5, through December 31, 20x6 was purchased on December 15, 20X4. The premium paid totaled $24,000 and was recorded as pre- paid insurance. The operation purchased a used cash register costing $8,000 on December 18, 20X4 and recorded it in the equipment account. The cash register is expected to have a useful life of five years and a salvage value of $0. No depreciation has been recorded. Fees uses the straight-line method of depreciation. The employees were paid for their work through December 28. They worked 100 hours for the period of December 29-31, 20X5, and will be paid on January 10, 20X6. The average hourly wage is $8.00. The related payroll taxes are 10 percent of the wages 2· 3. 4. Sales for the year totaled $1,000,000. The allowance for doubtful accounts has a Decem- ber 31 balance of $1,500 prior to write-off of two accounts totaling $300. The allowance should be adjusted at year-end to ¾ percent of sales for the year.

Problem13
Answer #1

13)
1)insurance per month=24000/24=1000
insurance expense(db)1000*12=12000
prepaid insurance(Cr)12000

2)depreciaiton=(price-salvage)/years
=(8000-0)/5=1600
depreciation equipment(db)1600
accumulated depreciation-equipment(cr)1600

3)salary=100*8=800
payrol tax=10%*800-80
salary expense(db)800
payroll tax expnese(db)80
salaray payable(cr)800
payroll tax payable(Cr)80

4)allowance=0.75%*1000000=7500
bad debt expense(db)300
allowance for doutbful accounts(cr)300

already we have balance of 1500 credit and 300 extra and we need total to be 7500
remaining=7500-1500-300=5700
bad debt expense(db)5700
allowance for doutbful accounts(cr)5700

5)Electricity expense(db)320
electricity payable(cr)320